OpenStudy (neo92):

Solve the following deferential equation... y^4+2y^3+3y^2+2y+y=0

OpenStudy (shubhamsrg):

differential ?

OpenStudy (neo92):

yep.. my book says so.. y^(4)+2y^(3)+3y^(2)+2y(1)+y=0

OpenStudy (neo92):

I think (4) is the order

OpenStudy (anonymous):

I'm guessing there's an error in how the question was written but, assuming f(y) = y^(4)+2y^(3)+3y^(2)+2y(1)+y, d[f(y)]/dy = 4y^3 + 6y^2 + 6y + 2

OpenStudy (anonymous):

Might y^(4) mean the fourth derivative of y? Do you have any boundary conditions supplied?

OpenStudy (neo92):

@TimSmit I too think so. But no boundry conditions given.

OpenStudy (anonymous):

That is odd...well then take the trivial solution y = 0 :)